As a publicity stunt, a late night talk show host drops a pumpkin from a rooftop that is 200 feet high. When will the pumpkin hit the ground, the nearest hundredth of a second?
Answer by Alan3354(69443)

As a publicity stunt, a late night talk show host drops a pumpkin from a rooftop that is 200 feet high. When will the pumpkin hit the ground, the nearest hundredth of a second?
---------
s = at^2/2 = 16t^2
t = sqrt(s/16) = sqrt(200/16) seconds